A 1.0 mm -diameter, 20 cm -long copper wire carries a 4.5 A current.

What is the potential difference between the ends of the wire?
Answer #1

Here ,

current in wire, I = 4.5 A

resistance in wire , R = p * L/A

R = 1.68 *10^-8 * 0.20/(3.141 * (0.5 *10^-3)^2)

R = 4.279 *10^-3 Ohm

potential difference = I * R

potential difference = 4.5 * 4.279 *10^-3

potential difference = 0.0193 V

the potential difference between the ends of wire is 0.0193 V